LASER SPECTROSCOPY OF NiBr: ROTATIONAL ANALYSIS OF THE ^{2}Π_{3/2} − X^{2}Π_{3/2} TRANSITION A. S-C. CHEUNG, \underline{J. W-H. LEUNG}, JINGHUA DAI and XIANGHUAI WANG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong.